OpenStudy (anonymous):

A woman drives 6 miles, accelerating uniformly from rest to 70 mph. How long does it take for her to reach 70 mph?

OpenStudy (anonymous):

here u=0 ,v=70 s=6 and we need 't' use the relation s=((u+v)t)/2 hence here t=2s/(u+v)=2*6/70 hr=6/35 hr